OpenStudy (faman39):

Morissa pays $9.99 per pound for dried blueberries and $7.99 per pound for dried cherries. Which of the following statements could be true? Morissa bought 0.7 pounds of blueberries and 1.3 pounds of cherries for a total of $15. Morissa spent $16.25 on blueberries and spent $13.62 on cherries for a total of 3.0 pounds of fruit. Morissa bought 2 pounds of blueberries and spent $14.99 on cherries for a total of $29.97. Morissa bought 2 pounds of fruit for a total of $18.

OpenStudy (anonymous):

so you should be able to plug values in and see if they're true so for the first one multiply the pounds given by the price of each and add them together to see if it equals the amount given

OpenStudy (anonymous):

(9.99*0.7)=6.993 + (7.99*1.3)=10.387 = 17.38

OpenStudy (anonymous):

9.99 per lb for blueberries 7.99 per lb for cherries now lets look at answer choice A. (9.99)(.7) = 6.993 = 6.99 (7.99)(1.3) = 10.387 = 10.39 10.39 + 6.99 = 17.38 (incorrect)--does not equal $ 15 now lets look at answer choice B. 1/9.99 = x/16.25 (cross multiply) 1/7.99 = x/13.62 (9.99)(x) = (1)(16.25) (7.99)(x) = (1)(13.62) 9.99x = 16.25 7.99x = 13.62 x = 16.25/9.99 x = 13.62/7.99 x = 1.6 lbs x = 1.7 lbs now add them : 1.6 + 1.7 = 3.3 lbs (incorrect) - does not equal 3 lbs now lets look at answer choice 3 9.99 x 2 = 19.98 cherries ....14.99 ---------------add total = 34.97 (incorrect) -- does not equal 29.97 now lets look at answer choice 4 b + c = 2 ---> b = 2 - c 9.99b + 7.99c = 18 now sub 2 - c in for b in the second equation 9.99b + 7.99c = 18 9.99(2 - c) + 7.99c = 18 19.98 - 9.99c + 7.99c = 18 - 2c = 18 - 19.98 -2c = -1.98 c = 1.98/2 c = .99 now sub .99 in for c in the 1st equation b + c = 2 b + .99 = 2 b = 2 - .99 b = 1.01 Morissa bought 1.01 lbs of blueberries at 9.99/lb = 10.09 .99 lbs of cherries at 7.99/lb = 7.91 for a total of 2 lbs of fruit for (10.09 + 7.91) = 18 (correct) answer 4 is the answer. Do you have any questions about what I just did ?